Ryan puts 2 nickels, 2 dimes, and 1 quarter in a bag. He then picks one coin. What is the expected value, in cents, of each pick in his experiment?
a.13 cents
b.11 cents
c.10 cents
d.9 cents

Answer by Edwin McCravy(20055)

Ryan puts 2 nickels, 2 dimes, and 1 quarter in a bag. He then picks one coin. What is the expected value, in cents, of each pick in his experiment?


Pick     P(pick)   Value of pick   Value×P(pick)=E(pick)
nickel   2/5=.4        $.05             $.05×.4 = $.02    
dime     2/5=.4        $.10             $.10×.4 = $.04
quarter  1/5=.2        $.25             $.20×.2 = $.05
--------------------------------------------------------
            1.0                                   $.11

Expectation = 11 cents.

11 cents is what Ryan will expect to average getting per 
experiment if he were to repeat this same experiment many 
times.

Edwin